What is the dot product of two vectors of magnitude 3 and 5 if the angle between them is 60Degree

Grade:11

1 Answers

Om
6 years ago
Dot product of two vectors is satisfied by function of costheta. So3 • 5 = 3×5cos60 =15*1\2 =7•5And for cross product use sintheta